# BNMusic: Blending Environmental Noises into Personalized Music

### v1

@@ Abstract
While being disturbed by environmental noises, the acoustic masking technique is a conventional way to reduce the annoyance in audio engineering that seeks to cover up the noises with other dominant yet less intrusive sounds. However, misalignment between the dominant sound and the noise—such as mismatched downbeats—often requires an excessive volume increase to achieve effective masking. Motivated by recent advances in cross-modal generation, in this work, we introduce an alternative method to acoustic masking, aiming to reduce the noticeability of environmental noises by blending them into personalized music generated based on user-provided text prompts. Following the paradigm of music generation using mel-spectrogram representations, we propose a Blending Noises into Personalized Music (BNMusic) framework with two key stages. The first stage synthesizes a complete piece of music in a mel-spectrogram representation that encapsulates the musical essence of the noise. In the second stage, we adaptively amplify the generated music segment to further reduce noise perception and enhance the blending effectiveness, while preserving auditory quality. Our experiments with comprehensive evaluations on MusicBench, EPIC-SOUNDS, and ESC-50 demonstrate the effectiveness of our framework, highlighting the ability to blend environmental noise with rhythmically aligned, adaptively amplified, and enjoyable music segments, minimizing the noticeability of the noise, thereby improving overall acoustic experiences.
@@ 1 Introduction
Our proposed task draws inspiration from the theory of auditory masking [6, 33, 26], a psychoacoustic phenomenon where the perception of one sound is reduced or eliminated by the presence of another sound. This effect is typically modeled in the frequency domain using psychoacoustic principles, which define a masking threshold—the minimum intensity below which a sound becomes inaudible to the human ear in the presence of a masker. When a new sound is introduced, it elevates the masking threshold not only at its own frequency but also across neighboring frequency bands, making weaker signals in those regions imperceptible. As shown in Fig. 1 (b), such thresholds are fundamental to many perceptual audio models, and here we use them to guide the blending of generated music with environmental noise. This task aims to diminish the perception of unwanted noise by introducing supplementary background sounds. In practice, fully masking complex environmental noise within a comfortable loudness range is rarely achievable. Instead, we aim for a more feasible and perceptually effective solution—blending. By generating music that is rhythmically aligned with the underlying noise, our approach enables partial masking while incorporating the residual noise components into the musical texture, thus reducing perceived annoyance without overwhelming the listener.
@@
To effectively blend generated music with environmental noise, it is essential that the music can perceptually mask the noise, thereby reducing its distracting impact. Achieving such blending requires that the masking effect remain strong even at relatively low overall loudness levels. To this end, the generated music should align closely with the noise in terms of rhythm or other structural properties, enabling it to integrate naturally without relying on excessive loudness. However, while recent advances in music generation—particularly those based on mel-spectrogram representations in the frequency domain [5, 34, 18, 13, 38, 39, 42, 16]-have demonstrated remarkable progress, most existing models are trained to generate music from clean, structured inputs such as text prompts or orderly music excerpts. These models struggle when conditioned on noisy and unstructured inputs, as they lack the ability to retain relevant acoustic attributes from chaotic signals. Yet they also provide a natural foundation for leveraging masking principles, which are inherently frequency-dependent. Building on this emerging paradigm, we explore generating music in the frequency domain whose structure, rhythmic patterns, and statistical properties naturally align with those of the noise, enabling seamless auditory blending, which in turn diminishes the listeners’ awareness of the underlying noise.
In this paper, we propose a novel approach Blending Noises into Personalized Music (BNMusic) that uses adaptive loudness-amplified music to blend with background noise. This approach makes the noise less noticeable while maintaining a balanced overall volume, thereby enhancing the acoustic environment effectively. Our method achieves seamless alignment and integration of generated music with ambient noise through a two-stage process. In the first stage, we generate music that blends rhythmically with the noise by applying a two-step outpainting and inpainting process on the noise mel-spectrogram representation to separately reconstruct a complete piece of music in the frequency domain. Outpainting is used to generate music around the peak energy region of the noise spectrogram, followed by inpainting to reconstruct music within this high-energy zone. This sequence diffuses musical qualities from the noise’s core to the surrounding areas and back into the core, effectively integrating the noise’s inherent characteristics into a cohesive musical composition. Through this process, the method finds an optimal balance between maintaining the noise’s relevance and ensuring the resulting music remains the most pleasing to the ear. In the second stage, we dynamically amplify each music segment’s loudness to enhance the masking effect in the core area without increasing the overall volume excessively. This reduces the perceptibility of noise where its energy is most concentrated while facilitating a more organic integration of the residual components with the music, thereby enhancing the overall auditory experience. In this targeted manner, our method achieves a robust and effective blending, integrating the generated music with the noise so that the noise becomes unnoticeable to the listener with the total volume maintaining pleasing.
@@
To assess the efficacy of our approach in blending with environmental noise and creating a more pleasant auditory experience, we conduct comprehensive objective and subjective evaluations. We use EPIC-SOUNDS[9] and ESC-50[30] as noise sources, which together cover a wide range of real-world environmental sounds, including various actions, objects, and acoustic scenes, spanning a broad frequency range from 100 Hz to 10,000 Hz. Results show that our complete method achieves the best performance on MusicBench [18]. In summary, we present three major contributions:
@@ 2 Related work
Conventional acoustic methods. The conventional acoustic methods of enhancing hearing environments have evolved significantly throughout these years, driven by the demand for better auditory experiences. Early efforts focused on employing physical barriers and materials to block out noise. The introduction of active noise canceling (ANC) [11] marked a major breakthrough, which utilizes microphones and electronic circuitry to produce anti-phase sound waves that counteract noise [35]. Researchers have also studied auditory masking for years [6, 33], which describes how a louder sound can reduce or eliminate the perception of a quieter one occurring simultaneously. Building psychoacoustic models to simulate the human perception of audio [26, 22, 20] and quantifying how much louder a masker needs to be at different frequencies to effectively mask a target sound. Studies have applied auditory masking principles to reduce the perceptual impact of noise, designing comfortable soundscapes using chord progressions and melodies that align with the peak frequencies of disruptive noises [21] and investigating the use of natural sounds to mask traffic noise in urban environments [41]. Or utilizing audio masking for audio watermarking [37]. Inspired by these masking-based techniques, we propose the noise blending task that uses the auditory masking effect to reduce the perceptual impact of noise on acoustic environments.
@@
Generative models for music. Generative models have recently achieved remarkable success in both vision and audio synthesis [5, 12, 34, 18, 28, 4]. In image generation, diffusion models [8, 31] have proven especially effective by gradually denoising in the latent space of a pre-trained network, leveraging its structure to produce high-quality outputs. Inspired by this success, diffusion-based techniques have also been extended to Text-to-Music tasks [12, 34, 18], significantly improving both generation quality and efficiency. Contemporary audio generation models generally fall into two categories: those that generate audio directly in the waveform domain, and those that first generate time-frequency representations (e.g., mel-spectrograms) and convert them to waveforms using a vocoder. MusicGen [3] exemplifies the former, synthesizing music through multiple streams of discrete audio tokens. In contrast, Riffusion [5] follows the latter approach, generating mel-spectrograms via a fine-tuned Stable Diffusion [31] model and converting them into audio. AudioLDM [12] builds on latent-space modeling and CLAP (Contrastive Language-Audio Pretraining) to generate audio from text. Mustango [18] introduces MuNet for fine-grained control, achieving competitive performance even with limited training data. AudioLDM2 [13] proposes a unified "language of audio" (LOA) representation, using GPT-2 to bridge multiple modalities and guiding generation with a latent diffusion model. Our work builds on these foundations, aiming to generate music that blends harmoniously with environmental noise using diffusion-based methods. Meanwhile, complementary efforts have focused on accelerating generation for real-time applications [24, 32, 23, 2], which may enhance the practical deployment of our method in future iterations.
@@
Generation controlling techniques. Recent years have witnessed significant progress in audio editing, with a growing number of works exploring the potential of generative models in transforming and manipulating sound [27, 38, 19, 1, 17, 39, 42, 16]. Many approaches adopt a vision-inspired paradigm by converting audio into 2D representations such as mel-spectrograms, enabling the application of powerful image editing techniques to audio [5, 34, 18, 13, 38, 39, 42, 16]. Following the advances in image generative models like Stable Diffusion [31], audio editing has expanded to include tasks such as spectrogram inpainting, style transfer, and attribute control. For instance, Audit [38] applied latent diffusion models (LDMs) to edit audio in a controllable manner. Prior audio inpainting methods [19, 1, 17] typically focus on interpolating the gap between two audio clips. In contrast, more recent works [16, 42, 7, 14] leverage spectrogram-based image diffusion to manipulate high-level features such as genre, instrument, or mood. Inspired by inpainting tasks in vision and their success in image composition and context-aware filling [15, 40], we propose a new application scenario in which spectrogram inpainting and outpainting are used to mask environmental noise through musical blending. Built upon Riffusion [5], our method adapts image-style spectrogram generation to synthesize rhythmically aligned, stylistically coherent music that perceptually reduces the annoyance of background noise. Unlike traditional gap-filling methods, our approach treats the noise-occupied spectrogram as a canvas to be expanded and enhanced, thus offering a novel direction in content-aware audio editing.
@@ 3 BNMusic framework: Blending Noises into personalized Music
As the input noise $A_{\text{Noise}}\in\mathbb{R}^{t\times f_{s}}$, where $t$ is the length of audio in seconds and $f_{s}$ is the sampling rate (i.e., number of samples per second), is given to the system, we convert it from an audio signal into a mel-spectrogram $\mathbf{S}_{\text{Noise}}=\text{Mel}|\text{STFT}(A_{\text{Noise}})|\in\mathbb{R}^{W\times H}$, where Mel stands for the Mel-filtering process and STFT means Short Time Fourier Transform. Through this conversion, the input one-dimensional signal $A_{\text{Noise}}$ can be represented in a two-dimensional matrix - mel-spectrogram $\mathbf{S}_{\text{Noise}}$. Once the mel-spectrogram is obtained, its amplitude values are mapped to grayscale pixel intensities in the range $[0,255]$. In this mapping, lower pixel values represent louder regions, while higher values indicate quieter ones. This process yields a detailed grayscale mel-spectrogram plot, denoted as $\mathbf{x}_{\text{Noise}}\in[0,255]^{W\times H\times 1}$, where the last dimension represents a single grayscale channel. With such a more perceptually relevant representation of the signal, the evolution of noise frequency content over time becomes easier to analyze and extract. After obtaining $\mathbf{x}_{\text{Noise}}$, we then apply a binary mask $\mathbf{M}\in\left\{0,1\right\}^{W\times H}$ that masks out the area represent the higher energy part of the noise. The masked image can be calculated as $\widetilde{\mathbf{x}}_{\text{Noise}}=\mathbf{x}_{\text{Noise}}\odot\mathbf{M}\in[0,255]^{W\times H\times 1}$, where $\odot$ represents element-wise multiplication.
@@
We adopt a two-step outpainting and inpainting process to preserve the musical essence (e.g., the rhythmic patterns) from the input noise. This design facilitates the blending of the noise into music while maintaining harmony. First, the mask $\mathbf{M}$ isolates the core noise region in the image $\widetilde{\mathbf{x}}_{\text{Noise}}$, dividing it into two parts. During the outpainting stage, the core noise area is preserved, and music is generated to fill the surrounding space, allowing the core information to diffuse outward. The masked mel-spectrogram plot $\widetilde{\mathbf{x}}_{\text{Noise}}$ and the text prompt $C_{\text{text}}$ are encoded into latent representations, then fed into the LDM [31] from Riffusion [5] for music generation, which is a modified version of Stable Diffusion-v1-5 fine-tuned for generating music’s mel-spectrogram plot. Given a noisy latent representation $\mathbf{z}_{t}$ at timestep $t$, the model predicts the added noise $\boldsymbol{\epsilon}_{\theta}$ using a U-Net conditioned on both the corrupted mel-spectrogram $\widetilde{\mathbf{x}}_{\text{Noise}}$ and the associated text prompt $C_{\text{text}}$. Using the predicted noise $\boldsymbol{\epsilon}_{\theta}$, the posterior distribution of the previous latent state $\mathbf{z}_{t-1}$ is computed as:
@@
After a second round of inpainting on the inversely masked area on $\widetilde{\mathbf{x}}_{\text{Mid}}$, we obtain an image $\mathbf{x}_{\text{Music}}$, an image representing the complete musical content. This iterative process extracts rhythmic patterns from the noise and integrates them into the music, replacing the noisy elements. The second round of inpainting is crucial for refining the composition, ensuring that the information diffused to the surrounding areas is reintegrated into the core, ultimately reconstructing a piece of music that preserves the rhythmic characteristics of the original noise while eliminating distracting parts.
@@
To improve blending, we amplify the generated music to enhance its masking effect on the core area of the noise. The music incorporates rhythmic patterns extracted from the noise’s high-energy core area, making it easier to blend with the noise. By applying minimal loudness amplification, we effectively mask the core area’s mel-frequencies, making the noise less noticeable. To calculate the masking threshold, we first compute the noise’s magnitude spectrogram using the Short-Time Fourier Transform (STFT), which extracts the magnitude of the complex STFT coefficients. The resulting matrix, $\hat{\mathbf{S}}_{\text{Noise}}$, is real-valued and belongs to $\mathbb{R}^{\hat{W}\times H}$, where $\hat{W}$ is the number of frequency bins and $H$ is the number of time frames. Based on previous research [26], the minimum signal-to-mask ratio (SMR) for Tune-Masking-Noise cases is typically between 21–28 dB. Using the minimum value of 21 dB, we derive the threshold matrix as:

### v2

@@ Abstract
While being disturbed by environmental noises, the acoustic masking technique is a conventional way to reduce the annoyance in audio engineering that seeks to cover up the noises with other dominant yet less intrusive sounds. However, misalignment between the dominant sound and the noise—such as mismatched downbeats—often requires an excessive volume increase to achieve effective masking. Motivated by recent advances in cross-modal generation, in this work, we introduce an alternative method to acoustic masking, aiming to reduce the noticeability of environmental noises by blending them into personalized music generated based on user-provided text prompts. Following the paradigm of music generation using mel-spectrogram representations, we propose a Blending Noises into Personalized Music (BNMusic) framework with two key stages. The first stage synthesizes a complete piece of music in a mel-spectrogram representation that encapsulates the musical essence of the noise. In the second stage, we adaptively amplify the generated music segment to further reduce noise perception and enhance the blending effectiveness, while preserving auditory quality. Our experiments with comprehensive evaluations on MusicBench, EPIC-SOUNDS, and ESC-50 demonstrate the effectiveness of our framework, highlighting the ability to blend environmental noise with rhythmically aligned, adaptively amplified, and enjoyable music segments, minimizing the noticeability of the noise, thereby improving overall acoustic experiences. Project page: https://d-fas.github.io/BNMusic_page/.
@@ 1 Introduction
Our proposed task draws inspiration from the theory of auditory masking [7, 35, 28], a psychoacoustic phenomenon where the perception of one sound is reduced or eliminated by the presence of another sound. This effect is typically modeled in the frequency domain using psychoacoustic principles, which define a masking threshold—the minimum intensity below which a sound becomes inaudible to the human ear in the presence of a masker [5]. When a new sound is introduced, it elevates the masking threshold not only at its own frequency but also across neighboring frequency bands, making weaker signals in those regions imperceptible [22]. As shown in Fig. 1 (b), such thresholds are fundamental to many perceptual audio models, and here we use them to guide the blending of generated music with environmental noise. This task aims to diminish the perception of unwanted noise by introducing supplementary background musical sounds. In practice, fully masking complex environmental noise within a comfortable loudness range is rarely achievable. Instead, we aim for a more feasible and perceptually effective solution through blending. By generating music that is rhythmically aligned with the underlying noise, our approach enables partial masking while incorporating the residual noise components into the musical texture, thus reducing perceived annoyance without overwhelming the listener.
@@
To effectively blend generated music with environmental noise, it is essential that the music can perceptually mask the noise, thereby reducing its distracting impact. Achieving such blending requires that the masking effect remain strong even at relatively low overall loudness levels. To this end, the generated music should align closely with the noise in terms of rhythm or other structural properties, enabling it to integrate naturally without relying on excessive loudness. However, while recent advances in music generation, particularly those based on mel-spectrogram representations in the frequency domain [6, 36, 19, 14, 40, 41, 44, 17], have demonstrated remarkable progress, most existing models are trained to generate music from clean, structured inputs such as text prompts or orderly music excerpts. These models struggle when conditioned on noisy and unstructured inputs, as they lack the ability to retain relevant acoustic attributes from chaotic signals. Yet they also provide a natural foundation for leveraging masking principles, which are inherently frequency-dependent. Building on this emerging paradigm, we explore generating music in the frequency domain whose structure, rhythmic patterns, and statistical properties naturally align with those of the noise, enabling seamless auditory blending, thus in turn diminishes the listeners’ awareness of the underlying noise.
In this paper, we propose a novel approach Blending Noises into Personalized Music (BNMusic) that uses adaptive loudness-amplified music to blend with background noise. This approach makes the noise less noticeable while maintaining a balanced overall volume, thereby enhancing the acoustic environment effectively. Our method is designed as a two-stage process that explicitly targets the high-energy regions of noise, which are the most perceptually salient and thus the hardest to mask. In the first stage, we condition on these regions and apply a two-step outpainting–inpainting process on the noise mel-spectrogram to generate music that rhythmically and spectrally aligns with the noise. Outpainting extends musical patterns around the dominant noise zones, while inpainting reconstructs coherent content within them, ensuring that the resulting composition inherits the noise’s most prominent frequency structures in a natural way. This design effectively transforms disruptive noise components into elements that can be musically integrated. In the second stage, we apply adaptive amplification targeted to these same regions. Because the generated music already shares temporal–spectral characteristics with the noise, only modest gain is required to strengthen the masking effect, avoiding excessive loudness while maintaining musical balance. Together, the two stages form a tightly coupled system: Stage 1 embeds the noise into the music in a perceptually aligned manner, and Stage 2 consolidates this alignment by enhancing masking where it matters most. As a result, BNMusic achieves robust blending that suppresses the perceptual salience of noise while preserving coherence and listening comfort.
@@
To assess the efficacy of our approach in blending with environmental noise and creating a more pleasant auditory experience, we conduct comprehensive objective and subjective evaluations. We use EPIC-SOUNDS[10] and ESC-50[32] as noise sources, which together cover a wide range of real-world environmental sounds, including various actions, objects, and acoustic scenes, spanning a broad frequency range from 100 Hz to 10,000 Hz. Results show that our complete method achieves the best performance on MusicBench [19].
In summary, we present three major contributions:
@@ 2.1 Conventional acoustic methods
The conventional acoustic methods of enhancing hearing environments have evolved significantly throughout these years, driven by the demand for better auditory experiences. Early efforts focused on employing physical barriers and materials to block out noise. The introduction of active noise canceling (ANC) [12] marked a major breakthrough, which utilizes microphones and electronic circuitry to produce anti-phase sound waves that counteract noise [37]. Researchers have also studied auditory masking for years [7, 35], which describes how a louder sound can reduce or eliminate the perception of a quieter one occurring simultaneously. Building psychoacoustic models to simulate the human perception of audio [28, 24, 21] and quantifying how much louder a masker needs to be at different frequencies to effectively mask a target sound. Studies have applied auditory masking principles to reduce the perceptual impact of noise, designing comfortable soundscapes using chord progressions and melodies that align with the peak frequencies of disruptive noises [23] and investigating the use of natural sounds to mask traffic noise in urban environments [43]. Or utilizing audio masking for audio watermarking [39]. Inspired by these masking-based techniques, we propose the noise blending task that uses the auditory masking effect to reduce the perceptual impact of noise on acoustic environments.
2.2 Generative models for music
@@
Generative models have recently achieved remarkable success in both vision and audio synthesis [6, 13, 36, 19, 30, 4, 45]. In image generation, diffusion models [9, 33] have proven especially effective by gradually denoising in the latent space of a pre-trained network, leveraging its structure to produce high-quality outputs. Inspired by this success, diffusion-based techniques have also been extended to Text-to-Music tasks [13, 36, 19, 45], significantly improving both generation quality and efficiency. Contemporary audio generation models generally fall into two categories: those that generate audio directly in the waveform domain, and those that first generate time-frequency representations (e.g., mel-spectrograms) and convert them to waveforms using a vocoder. MusicGen [3] exemplifies the former, synthesizing music through multiple streams of discrete audio tokens. In contrast, Riffusion [6] follows the latter approach, generating mel-spectrograms via a fine-tuned Stable Diffusion [33] model and converting them into audio. AudioLDM [13] builds on latent-space modeling and CLAP (Contrastive Language-Audio Pretraining) to generate audio from text. Mustango [19] introduces MuNet for fine-grained control, achieving competitive performance even with limited training data. AudioLDM2 [14] proposes a unified "language of audio" (LOA) representation, using GPT-2 to bridge multiple modalities and guiding generation with a latent diffusion model. Our work builds on these foundations, aiming to generate music that blends harmoniously with environmental noise using diffusion-based methods. Meanwhile, complementary efforts have focused on accelerating generation for real-time applications [26, 34, 25, 2], which may enhance the practical deployment of our method in future iterations.
2.3 Generation controlling techniques
@@
Recent years have witnessed significant progress in audio editing, with a growing number of works exploring the potential of generative models in transforming and manipulating sound [29, 40, 20, 1, 18, 41, 44, 17]. Many approaches adopt a vision-inspired paradigm by converting audio into 2D representations such as mel-spectrograms, enabling the application of powerful image editing techniques to audio [6, 36, 19, 14, 40, 41, 44, 17]. Following the advances in image generative models like Stable Diffusion [33], audio editing has expanded to include tasks such as spectrogram inpainting, style transfer, and attribute control. For instance, Audit [40] applied latent diffusion models (LDMs) to edit audio in a controllable manner. Prior audio inpainting methods [20, 1, 18] typically focus on interpolating the gap between two audio clips. In contrast, more recent works [17, 44, 8, 15] leverage spectrogram-based image diffusion to manipulate high-level features such as genre, instrument, or mood. Inspired by inpainting tasks in vision and their success in image composition and context-aware filling [16, 42], we propose a new application scenario in which spectrogram inpainting and outpainting are used to mask environmental noise through musical blending. Built upon Riffusion [6], our method adapts image-style spectrogram generation to synthesize rhythmically aligned, stylistically coherent music that perceptually reduces the annoyance of background noise. Unlike traditional gap-filling methods, our approach treats the noise-occupied spectrogram as a canvas to be expanded and enhanced, thus offering a novel direction in content-aware audio editing.
@@ 3 BNMusic framework: Blending Noises into personalized Music
As the input noise $A_{\text{Noise}}\in\mathbb{R}^{t\times f_{s}}$, where $t$ is the length of audio in seconds and $f_{s}$ is the sampling rate (i.e., number of samples per second), is given to the system, we convert it from an audio signal into a mel-spectrogram $\mathbf{S}_{\text{Noise}}=\text{Mel}|\text{STFT}(A_{\text{Noise}})|\in\mathbb{R}^{W\times H}$, where Mel stands for the Mel-filtering process and STFT means Short Time Fourier Transform. Through this conversion, the input one-dimensional signal $A_{\text{Noise}}$ can be represented in a two-dimensional matrix, namely the mel-spectrogram $\mathbf{S}_{\text{Noise}}$. Once the mel-spectrogram is obtained, its amplitude values are mapped to grayscale pixel intensities in the range $[0,255]$. In this mapping, lower pixel values represent louder regions, while higher values indicate quieter ones. This process yields a detailed grayscale mel-spectrogram plot, denoted as $\mathbf{x}_{\text{Noise}}\in[0,255]^{W\times H\times 1}$, where the last dimension represents a single grayscale channel. With such a more perceptually relevant representation of the signal, the evolution of noise frequency content over time becomes easier to analyze and extract. After obtaining $\mathbf{x}_{\text{Noise}}$, we then apply a binary mask $\mathbf{M}\in\left\{0,1\right\}^{W\times H}$ that masks out the area representing the higher energy part of the noise. Since frequencies with higher local energy are typically more perceptually salient and more likely to interfere with auditory perception, isolating these regions is crucial for subsequent blending. The masked image can be calculated as $\widetilde{\mathbf{x}}_{\text{Noise}}=\mathbf{x}_{\text{Noise}}\odot\mathbf{M}\in[0,255]^{W\times H\times 1}$, where $\odot$ represents element-wise multiplication.
@@
We adopt a two-step outpainting and inpainting process to guide the generation of music around and within the high-energy regions of the noise spectrogram. This design enables the model to capture and preserve the noise’s intrinsic rhythmic cues while diffusing them into a coherent musical structure, thereby ensuring that the generated music naturally aligns with the most perceptually salient parts of the noise. First, the mask $\mathbf{M}$ isolates the core noise region in the image $\widetilde{\mathbf{x}}_{\text{Noise}}$, dividing it into two parts. During the outpainting stage, the core noise area is preserved, and music is generated to fill the surrounding space, allowing the core information to diffuse outward. The masked mel-spectrogram plot $\widetilde{\mathbf{x}}_{\text{Noise}}$ and the text prompt $C_{\text{text}}$ are encoded into latent representations, then fed into the LDM [33] from Riffusion [6] for music generation, which is a modified version of Stable Diffusion-v1-5 fine-tuned for generating music’s mel-spectrogram plot. Given a noisy latent representation $\mathbf{z}_{t}$ at timestep $t$, the model predicts the added noise $\boldsymbol{\epsilon}_{\theta}$ using a U-Net conditioned on both the corrupted mel-spectrogram $\widetilde{\mathbf{x}}_{\text{Noise}}$ and the associated text prompt $C_{\text{text}}$. Using the predicted noise $\boldsymbol{\epsilon}_{\theta}$, the posterior distribution of the previous latent state $\mathbf{z}_{t-1}$ is computed as:
@@
After a second round of inpainting on the inversely masked area on $\widetilde{\mathbf{x}}_{\text{Mid}}$, we obtain an image $\mathbf{x}_{\text{Music}}$, an image representing the complete musical content. This iterative process extracts rhythmic patterns from the noise and integrates them into the music, replacing noisy elements while eliminating the most distracting parts. The second round of inpainting is crucial for refining the composition, ensuring that the diffused information is reintegrated into the core, ultimately reconstructing a piece of music that not only preserves the rhythmic characteristics of the original noise but also aligns with its dominant perceptual features, thereby laying a solid foundation for effective blending in the subsequent stage.
@@
To further enhance blending, we adaptively applies minimal loudness amplification, exploiting the music’s alignment with the noise’s high-energy regions to achieve effective auditory masking with modest gain, thereby reducing perceptual intrusiveness without disrupting the overall acoustic balance. To calculate the masking threshold, we first compute the noise’s magnitude spectrogram using the Short-Time Fourier Transform (STFT), which extracts the magnitude of the complex STFT coefficients. The resulting matrix, $\hat{\mathbf{S}}_{\text{Noise}}$, is real-valued and belongs to $\mathbb{R}^{\hat{W}\times H}$, where $\hat{W}$ is the number of frequency bins and $H$ is the number of time frames. Based on previous research [28], the minimum signal-to-mask ratio (SMR) for Tune-Masking-Noise cases is typically between 21–28 dB. Using the minimum value of 21 dB, we derive the threshold matrix as: